A wheel has a radius of 3 inches. About how far will it travel in 30 revolutions?
Answer by stanbon(75887)

A wheel has a radius of 3 inches. About how far will it travel in 30 revolutions?
---
circumference = 2*pi*r = 6pi
-----
Distance = 30(6pi) = 180pi is approximately 565.49 inches
===============================
Cheers,
Stan H.